OpenStudy (anonymous):

Solve for x: \log\ (x^(3)) = (\log\ x)^(2) Note, there are 2 solutions, A and B, where A < B.

OpenStudy (asnaseer):

try using the substitution:\[y=\log{x}\]and solve for y first.

OpenStudy (anonymous):

im still confused

OpenStudy (asnaseer):

on what part?

OpenStudy (anonymous):

the substitution

OpenStudy (asnaseer):

what do you get when you use that substitution?

OpenStudy (asnaseer):

what is the resulting equation in y?

OpenStudy (asnaseer):

do you know the various log rules?

OpenStudy (asnaseer):

e.g.:\[\log{x^n}=n\log{x}\]

OpenStudy (jdoe0001):

hmm, I read the above as \(\bf log(x^3) = (logx)^2 \)

OpenStudy (jdoe0001):

or \(\bf log(x^3) = (log(x))^2 \)

OpenStudy (anonymous):

the second one

OpenStudy (asnaseer):

yes - that is what I am taking it as

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

k thnks

OpenStudy (asnaseer):

so you understand how to proceed now?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

great! :)